OpenStudy (anonymous):

what is (9✖8 ➕eight and 2 thirds ? plz help me !

OpenStudy (anonymous):

9 x 82/3 multiply 8 by 3 add to 2 to get26/3 9 x 3 = 27/3 26/3 + 27/3 = 53/3

OpenStudy (anonymous):